OpenStudy (anonymous):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many bacteria are there after 8 hours? Round to the nearest integer and do not include units in your answer. B(t) = 4 e0.8t

OpenStudy (anonymous):

is it 2407????

OpenStudy (anonymous):

can you write B(t) out?

OpenStudy (anonymous):

bacteria(time)

OpenStudy (anonymous):

is that what u ment?

OpenStudy (anonymous):

No, is the function B(t)= 4*e*0.8*t or B(t)=4^(e*0.8*t)

OpenStudy (anonymous):

the first one i got it thanks